OpenStudy (anonymous):

can someone explain how sec(-pi/4) = sqrt2? I have never encountered a negative angle like this roflmao

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

Are you sure it has a negative?

OpenStudy (jdoe0001):

\(\bf sec\left(-\frac{\pi}{4}\right) = \sqrt{2}\\\quad \\ -----------------------\\ \textit{keep in mind that}\\ sec(\theta) = \cfrac{1}{cos(\theta)} \implies sec\left(-\frac{\pi}{4}\right) = \cfrac{1}{cos\left(-\frac{\pi}{4}\right)}\\ cos(-\theta) = cos(\theta) \implies \cfrac{1}{cos\left(-\frac{\pi}{4}\right)} \implies \cfrac{1}{cos\left(\frac{\pi}{4}\right)}\\\quad \\ -----------------------\\ sec\left(-\frac{\pi}{4}\right) = \sqrt{2} \implies \cfrac{1}{cos\left(\frac{\pi}{4}\right)} =\sqrt{2} \implies \cfrac{1}{\sqrt{2}} = cos\left(\frac{\pi}{4}\right)\)

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{\sqrt{2}} \implies \cfrac{1}{\sqrt{2}} \times \cfrac{\sqrt{2}}{\sqrt{2}} \implies \cfrac{\sqrt{2}}{2}\\ ------------------------\\ \cfrac{\sqrt{2}}{2}= cos\left(\frac{\pi}{4}\right)\) and you can check your Unit Circle for that

OpenStudy (jdoe0001):